Suppose C is a 3 x 3 matrix such that det (C) = 4. Find det ( C + C)
1
Expert's answer
2020-10-06T17:40:42-0400

Let the matrix be

"C = \begin{pmatrix}
 a & b & c\\
 d & e & f\\
g & h & i
\end{pmatrix}" , the determinant is "\det(C) = a(ei-hf)+ b(fg-di)+c(dh-ge) ."

"\det (C+C) = \det(2C) = \begin{vmatrix}
 2a & 2b & 2c\\
 2 d & 2e & 2f\\
2g & 2h & 2i
\end{vmatrix} = 2a(2e\cdot2i-2h\cdot2f)+ 2b(2f\cdot2g-2d\cdot2i)+2c(2d\cdot2h-2g\cdot2e) = 2^3\cdot\big(a(ei-hf)+ b(fg-di)+c(dh-ge)\big) = 8\cdot\det(C) = 32."